OpenStudy (anonymous):

A piece of paper with an area of 60 square inches is divided into two pieces so that the area of one is 2/3 the are of the other. The area (in square inches) of one of the pieces is (A) 15 (B) 20 (C) 24 (D) 30 (E) 45 Im guessing it's E...

OpenStudy (raffle_snaffle):

Why do you think it's E?

OpenStudy (anonymous):

Because 30 isn't the answer so anything before that can't be the answer either.

OpenStudy (anonymous):

@raffle_snaffle

OpenStudy (mertsj):

Let x be area of first piece 2/3 x = area of second piece x+2/3 x = 60

OpenStudy (anonymous):

I got 40/1 for the area of the second piece.

OpenStudy (mertsj):

How did you get that?

OpenStudy (mertsj):

What is x + 2/3 x?

OpenStudy (anonymous):

100...?

OpenStudy (mertsj):

1x + 2/3 x = 5/3 x

OpenStudy (anonymous):

Oh....

OpenStudy (anonymous):

x=60 right?

OpenStudy (mertsj):

I guess you don't know how to solve equations?

OpenStudy (anonymous):

not really c:

OpenStudy (mertsj):

If you have an equation like this: \[\frac{3}{4}x=12\]

OpenStudy (mertsj):

Solve it by multiplying both sides by the reciprocal of the coefficient of x

OpenStudy (anonymous):

huuh...?

OpenStudy (mertsj):

So you don't know what reciprocals are?

OpenStudy (anonymous):

I know what reciprocals are, but im not sure what coefficient means.

OpenStudy (mertsj):

in the expression 3x, 3 is the coefficient of x

OpenStudy (anonymous):

oh...I got 16/1

OpenStudy (mertsj):